A rectangular garden has an area of 40 yds^2. The length is two yards less than twice its width. Find the length and width of the garden. (Set up equations and solve)
Answer by nerdybill(7384)

A rectangular garden has an area of 40 yds^2. The length is two yards less than twice its width. Find the length and width of the garden. (Set up equations and solve)
Let w = width
then
2w-2 = length
.
w(2w-2) = 40
2w^2-2w = 40
w^2-w = 20
w^2-w-20=0
(w+4)(w-5) = 0
w = {-4, 5}
we can throw out the negative solution (extraneous) leaving
w = 5 yards
.
length:
2w-2 = 2(5)-2 = 10-2 = 8 yards